### 18 Mar 2024 v1

@@ 1 Introduction
Note that the dry mass is the most popular physical quantity that QPM can quantify in biology, but there exist other readouts of interest, as listed in Ref. BOE3_1757 , namely the optical volume, some phase/mass ratiometric quantities, sphericity/eccentricity indices, phase kurtosis and skewness. Also, more advanced tomography techniques can also measure cell volume and refractive index distribution in 3 dimensions.
@@ 2.1 Bright-field microscopy
Note that all the equations above assume a scalar electric field. This scalar approximation in optics is usually employed and valid in two cases: (i) with a linearly polarized illumination, where the scattered light is assumed to globally retain the incident polarisation, and (ii) with an unpolarized illumination, where the two components of the electric are likely to remain identical. However, in some cases, the scalar approximation is no longer valid, in particular when studying anisotropic particles or birefringent materials. The more general formalism involves two scalar fields, corresponding to the $x$ and $y$ components of the electric field, and Eqs. (3) and (4) becomes more generally:
@@ 2.2 Off-axis techniques
Off-axis techniques involve the interference of multiple waves, which create fringes on the camera sensor. These fringes are processed using a demodulation algorithm to retrieve the phase or wavefront of the light beam. This family includes DHM, CGM, and DPM.
@@ 2.2.2 Cross-grating wavefront microscopy
QLSI is not the only technique that is based on the positioning of an optical element at the vicinity of a camera, as a means to gain more information than $|\mathbf{E}|^{2}$. In particular, a microlens arrays JRS17_S573 ; AOP10_512 or a diffusing plate SM6_2100737 can also be placed at the vicinity of a camera. Diffuser-based phase sensing and imaging (DIPSI) uses a simple thin diffuser to retrieve the wavefront profile of a light beam SM6_2100737 , with a slightly reduced spatial resolution compared with QLSI. Shack-Hartmann (SH) wavefront imaging uses an array of microlenses to recover a wavefront profile JRS17_S573 ; OC222_81 . The spatial resolution is even lower than QLSI or DIPSI, making SH imaging rarely used in microscopy OL42_2122 . Light field microscopy (LFM) also uses an array of microlenses, but to retrieve another hidden information: the angular distribution of the light impinging at the sensor plane AOP10_512 ; BCJ16_397 . We propose to refer to these techniques as sensor-proximity techniques, i.e. techniques that place an optical element close to the camera sensor that redistributes the light intensity profile to reveal hidden information. Information theory says that if we gain information somewhere, we should also loose some information elsewhere. Indeed, with this technique family, this gain in information is accompanied by a loss in image definition.
@@ 2.3 Illumination-based differential phase contrast (DPC)
DPC is based on the sequential acquisition of a set of intensity images, using a standard bright field microscope, with various illumination conditions. For this purpose, a DPC microscope involves a large numerical aperture in illumination (typically 0.4, or even less), created by an extended source of light located in the front focal plane of the condenser (or far enough from the sample in the case there is no Köhler illumination) OE23_11394 . This extended source is uniformly switched on in the case of bright field microscopy (Fig. 2(a)). In the case of DPC, the light source is cut into two parts about its center, named left and right (Fig. 2(e)). Each half of the illumination gives rise to an asymmetric illumination. The two corresponding images, $I_{\mathrm{left}}$ and $I_{\mathrm{right}}$, are acquired. Then, the light source is split about the orthogonal axis to generate two other illuminations and two new images that we name $I_{\mathrm{top}}$ and $I_{\mathrm{bottom}}$. This set of 4 images are used to generate the DPC images defined by OE23_11394
@@
where $B$ is a uniform image, corresponding to the average intensity and $\ast$ denotes the convolution product. $H_{\mathrm{abs}}$ and $H_{\mathrm{pha}}$ are like point-spread-functions (PSFs) in intensity and phase. In this expression, only the last (phase) term depends on the illumination side (right or left, top or bottom). Here lies the interest of considering the image subtraction introduced above (Eqs. (10)), which become:
@@ 4.4.2 DPC accuracy: influence of Tikhonov regularization parameter $\alpha$
The other important degree of freedom in DPC is not experimental, but numerical. The Tikhonov regularization parameter $\epsilon$ is an artificial offset in the denominator of Eq. (16). It avoids division by small values to prevent a high noise level. The side effect of this artificial offset is a damping of the retrieved $\varphi$. Reference OE23_11394 mentions that a good compromise is $\epsilon=10^{-3}$. It is the value we used in all the simulations presented in Figs. 6 and 7. In Fig. 12, we compare images of neurons obtained with $\alpha=10^{-3}$ and $\epsilon=10^{-5}$. While the halo/shade-off effect is important for $\epsilon=10^{-3}$, it is well reduced for $\epsilon=10^{-5}$. This latter value is therefore more recommended, especially if quantitative dry mass measurements is the aim of the study. It, however, comes along a high noise level, as observed in Figs. 12d–f, which can be discarded by average multiples images, or using a camera with a large full-well-capacity.

### 22 Mar 2024 v2

@@ 1 Introduction
Note that the dry mass is the most popular physical quantity that QPM can quantify in biology, but there exist other readouts of interest, as listed in Ref. BOE3_1757 , namely the optical volume, some phase/mass ratiometric quantities, sphericity/eccentricity indices, phase kurtosis and skewness. Also, more advanced tomography techniques can also measure cell volume and refractive index distribution in 3 dimensions Lee_Hugonnet_Park_2022 ; Picazo-Bueno_Mico_2023 .
@@ 2.1 Bright-field microscopy
Note that all the equations above assume a scalar electric field. This scalar approximation in optics is usually employed and valid in two cases: (i) with a linearly polarized illumination, where the scattered light is assumed to globally retain the incident polarisation, and (ii) with an unpolarized illumination, where the two components of the electric field are likely to remain identical. However, in some cases, the scalar approximation is no longer valid, in particular when studying anisotropic particles or birefringent materials. The more general formalism involves two scalar fields, corresponding to the $x$ and $y$ components of the electric field, and Eqs. (3) and (4) becomes more generally:
@@ 2.2 Off-axis techniques
Off-axis techniques involve the interference of multiple waves, which create fringes on the camera sensor. These fringes are processed using a demodulation algorithm to retrieve the phase or wavefront of the light beam. This family includes DHM, CGM/QLSI, and DPM.
Unlike all the other techniques families described further on, off-axis techniques are based on the acquisition of a single image, containing all the information of intensity and phase. Information theory says that if we gain information somewhere compared with bright field microscopy (here the phase), we should also loose some information elsewhere. Indeed, with this technique family, this gain in information is always accompanied by a loss in image definition (the processed images have less pixels).
@@ 2.2.2 Cross-grating wavefront microscopy
QLSI is not the only technique that is based on the positioning of an optical element at the vicinity of a camera, as a means to gain more information than $|\mathbf{E}|^{2}$. In particular, a microlens arrays JRS17_S573 ; AOP10_512 or a diffusing plate SM6_2100737 can also be placed at the vicinity of a camera. Diffuser-based phase sensing and imaging (DIPSI) uses a simple thin diffuser to retrieve the wavefront profile of a light beam OL42_5117 ; SM6_2100737 , with a slightly reduced image definition (i.e. pixels number) compared with QLSI. Shack-Hartmann (SH) wavefront imaging uses an array of microlenses to recover a wavefront profile JRS17_S573 ; OC222_81 . The image definition is even lower than QLSI or DIPSI, making SH imaging rarely used in microscopy OL42_2122 . Light field microscopy (LFM) also uses an array of microlenses, but to retrieve another hidden information: the angular distribution of the light impinging at the sensor plane AOP10_512 ; BCJ16_397 . We propose to refer to these techniques as sensor-proximity techniques, i.e. techniques that place an optical element close to the camera sensor that redistributes the light intensity profile to reveal hidden information.
@@ 2.3 Illumination-based differential phase contrast (DPC)
In microscopy, the term DPC regroup different techniques that modify either the illumination or the detection angles to increase the contrast and retrieve quantitative information. In this paper, we focus on the most common implementation in QPM, that is DPC based on the sequential acquisition of a set of intensity images, using a standard bright field microscope, with various illumination conditions (illumination-based DPC). For this purpose, a DPC microscope involves a large numerical aperture in illumination (typically 0.25 to 0.4), created by an extended source of light located in the front focal plane of the condenser (or far enough from the sample in the case there is no Köhler illumination) OE23_11394 ; CB5_794 . This extended source is uniformly switched on in the case of bright field microscopy (Fig. 2(a)). In the case of DPC, the light source is cut into two parts about its center, named left and right (Fig. 2(e)). Each half of the illumination gives rise to an asymmetric illumination. The two corresponding images, $I_{\mathrm{left}}$ and $I_{\mathrm{right}}$, are acquired. Then, the light source is split about the orthogonal axis to generate two other illuminations and two new images that we name $I_{\mathrm{top}}$ and $I_{\mathrm{bottom}}$. This set of 4 images are used to generate the DPC images defined by OE23_11394
@@
where $B$ is the (uniform) intensity map at the image plane measured without object and $\ast$ denotes the convolution product. $H_{\mathrm{abs}}$ and $H_{\mathrm{pha}}$ are like point-spread-functions (PSFs) in intensity and phase. In this expression, considering an aberration-free system, only the last (phase) term depends on the illumination side (right or left, top or bottom). Here lies the interest of considering the image subtraction introduced above (Eqs. (10)), which become:
@@
The aim of the article is to compare microscopy techniques on the exact same microscopes. However, DPC has never been used with high-NA objectives, because the NA of the objective has to match the NA of the illumination. For this reason, in the DPC simulations, as an exception, we considered an objective NA of 0.4, instead of 1.3.
@@ 4.4.2 DPC accuracy: influence of Tikhonov regularization parameter $\alpha$
The other important degree of freedom in DPC is not experimental, but numerical. The Tikhonov regularization parameter $\epsilon$ is an artificial offset in the denominator of Eq. (16). It avoids division by small values to prevent a high noise level. The side effect of this artificial offset is a damping of the retrieved $\varphi$. Reference OE23_11394 mentions that a good compromise is $\epsilon=10^{-3}$. It is the value we used in all the simulations presented in Figs. 6 and 7. In Fig. 12, we compare images of neurons obtained with $\epsilon=10^{-3}$ and $\epsilon=10^{-5}$. While the halo/shade-off effect is important for $\epsilon=10^{-3}$, it is well reduced for $\epsilon=10^{-5}$. This latter value is therefore more recommended, especially if quantitative dry mass measurements is the aim of the study. It, however, comes along with a high noise level, as observed in Figs. 12d–f, which can be discarded by average multiples images, or using a camera with a large full-well-capacity. With $\epsilon=10^{-5}$, the halo/shade-off is reduced, but some non-uniformity remains in the background.
@@ 6 Summary
Finally, note that our numerical tool could also be used to compare 3D tomographic reconstructions provided by various QPM techniques. Indeed, IF-DDA can calculate the total field in any defocused plane with respect to the microscope focal plane, and for any illumination angle. With such data, 3D reconstructions can be easily studied.